An accountant works for TrailBlazer. On September 5th, he received an invoice totalling $32,500 from Whirlco with payment terms 3/10, 1.5/20, n/30. He sent in a partial payment on September 14th to reduce their balance by $12,500 and planned to pay off the remaining balance in time to receive the 1.5% cash discount.

a. What is the amount of the partial payment made on September 14th?

b. How much money would TrailBlazer lose if they don't get the second discount?
Answer #1
a
Amount of partial payment made on September 14th = 12500*(1-0.03)= $12125
b
Amount lost if they don't get the second discount = (32500-12500)*1.5%= $300